Circular No. 6240

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP) Z. Sekanina, Jet Propulsion Laboratory, writes: “The general scenario

for the bright jet’s evolution (IAUC 6223) remains plausible, with the sequence of high-resolution images taken between mid-August and early September (and now available on the World Wide Web) allowing one better to constrain the ejection times and particle velocities. The spin period is still indeterminate and is assumed to be 8 days. The best fit to the jet evolution is then achieved if the ejection from the equatorial source lasted from Aug. 17 through 19 (a quarter of the period) with velocities between 30 and 50 m/s. The position of the rotation axis is slightly revised to α = 40^{o}, δ = 0^{o} (obliquity 150^{o}). One may have to introduce a second source to simulate the innermost part of the jet at p.a. 290^{o}. The activity of this source, at a moderate (but poorly defined) ‘southern’ latitude, appears to have been of a very short duration, some time on Aug. 22, with ejection velocities of generally < 40 m/s.”

M. Kidger, Instituto di Astrofisica de Canarias, reports that this comet has been observed virtually every night in September with the 0.82-m IAC-80 Telescope at Teide Observatory (sequences of 300-s exposures in R): “The latest images show the development of a new and extremely bright jet. Since the decay of the spiral jet observed up to Sept. 6, the nuclear condensation had been perfectly circular, with no evidence of any important jet activity (through Sept. 25.88 UT). However, on Sept. 26.88, an elongation of the nuclear condensation was oriented east-west. By Sept. 27.88, this elongation had rotated anticlockwise by ∼ 45^{o} (to p.a. 35^{o}) and developed considerably into a jet similar to that observed in late August. On Sept. 28, a long bright ‘spiral’ arm extends clockwise from p.a. 323^{o} (where the jet is straight) to 46^{o}; the distance from the nucleus to the point in the jet where it starts to sweep in a clockwise direction is 4′′.5 on both Sept. 27 and 28. From this motion, we estimate a rotation period of ∼ 4.9 days. The new images may be viewed on the World Wide Web at

URL http://www.ll.iac.es/general/index.html.”

R. H. McNaught, Anglo-Australian Observatory, writes: “Images with the 1.0-m reflector on Sept. 29 UT show the comet to have a near-stellar nucleus with a high-intensity 3′′ jet in p.a. 320^{o}. The jet was only slightly fainter than the nucleus, which seems to lie on the southern tip of the inner coma/tail, the jet comprising the southwest boundary of the coma.”

1995 September 29 Daniel W. E. Green